# Hardware and Software Platform Inference

### v1

@@ 5.2 White-box Attacks
In the white-box setup, as explained in Section 4.1, we have access to the target model and its internal states. We can generate model-specific border inputs (HSPI-BI) or train a classifier on the output logits of the target model (HSPI-LD) to distinguish between different quantization levels or GPUs.
@@
GPUs FP32 BF16 FP16 MXINT8 FP8-E3 FP8-E4 INT8 RTX8000 vs H100 RTX8000 vs 2080Ti
Table 1: Table showing success in creating border images for different GPUs in white-box with an inference time batch size of 1.
@@
The HSPI-BI method always works to distinguish between all quantization levels in the white-box settings. Examples for border images can be found in Figure 4. The difference in logit values between quantization levels is visualized with the kernel density estimate in Figure 5. Extending to GPUs, the difference in logit bit distribution between RTX8000 and A100 can be seen in Figure 6. While for some GPUs the logit bit difference is quite significant, for others our tests show identical or too similar logit outputs, making it impossible to create border images. As can be seen in Table 1, for example, border images can be created between RTX8000 and H100, however not between RTX8000 and RTX2080Ti. Our tests also show that creating border images is simpler for larger batch sizes, but for consistency in Table 1, we focus on batch size 1. For instance, we found that RTX8000 vs H100 failed at batch size 1 but succeeded at 32 for FP8-E3 and FP8-E4. Section 6.1 discusses that further.
With HSPI-LD, we are able to distinguish between all quantization levels for the models ResNet18, ResNet50, VGG16, EfficientNet, DenseNet, and MobileNet-V2 with 100% accuracy when using an SVM with a set of 10 images’ logits per sample. Depending on the model, slightly different random creations of images performed better than others to most effectively capture the distribution. When comparing across GPUs, we are able to distinguish the A100 well from the RTX8000 and RTX2080Ti with an accuracy of $100\%$ for precisions FP32, BF16 and FP16. The difference in logits is also shown in Figure 6. However, we are not able to distinguish between the RTX8000 and RTX2080Ti, as there was no numerical difference between these two across all logits tested. In this way, we see the limitations to both methods in the case of two GPUs with similar architectures, where calculations are in the same EQC.
Figure 7: Example border request of DistillGPT2 generated by HPI-BI. When the border request is sent to the same model checkpoint deployed in FP16 and BF16, we observe different responses.
@@
We find that HSPI-BI does not work well with LLMs. With enough iterations, HSPI-BI can generate border requests that lead to different responses across quantization levels. Figure 7 illustrates an example of border requests for LLMs. However, we find that the PGD process is unstable due to the discontinuity in its projection. Moreover, limited GPU memory constrains our ability to scale experiments to larger LLMs and bigger request batch sizes. This is because the forward pass of the target model is part of the computation graph for gradient descent, requiring caching the intermediate activations of LLMs in memory for backward propagation.
@@
Figure 8: The difference of bit distribution between RTXA6000 and A100 (white-box HSPI-LD). We send the same 256 requests to QWen-2.5-3B deployed on RTXA6000 and A100 and compare the bit distribution of FP32 log probabilities generated by the model. Tokens and logits are sampled in the plot but the difference is still obvious. $\mathrm{t_{i}l_{j}}$ denotes the log probability of $i$-th token’s $j$-th logit.
On the contrary, HSPI-LD achieves remarkable success in both quantization experiments and actual GPU experiments. Table 2 shows the results of quantized LLaMA-3.1-8B, Phi-3.5-mini, and QWen-2.5-3B deployed in FP16, BF16, INT8-FD, and HQQ-4bit, with an accuracy over 99.5% for each model. When extending to actual GPUs, we first perform three initial experiments in Table 3 to verify that HSPI-LD can distinguish between different GPUs, arithmetic modes, and kernel implementations. We then conduct a comprehensive experiment in Table 4, mixing all these factors. We treat this experiment as a classification task with 18 unique labels. Remarkably, HSPI-LD achieves an overall accuracy of 83.9% (random guess accuracy = 5.6%).
@@ 5.3 Black-box Attacks
In the black-box setup, as explained in Section 4.1, we do not have access to or the knowledge of the underlying model. For both HSPI-BI and HSPI-LD, we presuppose the availability of a support set of models $\mathcal{F}_{s}$, with the understanding that the target model $F_{t}$ is excluded from this collection. Hence, we construct the border inputs (for HSPI-BI) or the classifier (for HSPI-LD) using $\mathcal{F}_{s}$, expecting that their functionality will transfer to the unseen model $F_{t}$.
@@
  • If border images are trained in batches, a larger batch size will improve the transferability across models. Figure 9(a) sweeps the batch size of border images trained on MobileNet-v3-Small and plots the resultant transferability. With the increasing batch sizes, the transferability is enhanced. However, we cannot further increase the batch size after 256 due to limited GPU memory.
  • Overfitting happens if the border images are trained with too many iterations. Figure 9(b) presents the transferability of the same group of border images trained with various number of iterations. The decreasing transferability at 800 and 1200 iterations indicates that the border images overfit the training model.
  • The border images trained on a more complex model tend to have better transferability. For example, in  Table 5, the border images trained on MobileNet-v2 has a higher F1-Score than those trained on ResNet18 and VGG16. However, again, more complex models require more GPU memory during training, which leads to limited scalability of HSPI-BI.
For HSPI-LD, we compare how well an SVM trained on one model performs on all other models found in  Table 5. We can see that the transferability in detecting BF16 and FP16 is particularly high, while for other quantization levels such as FP32 and MXINT8 the transferability is very low. Furthermore, ResNet18, followed by DenseNet perform best as models with high transferability to other models.
@@ 6 Discussion
6.1 Limitations and Black-box Failure Cases
@@
6.2 Software Supply Chains
The variations in the software supply chain can significantly influence the reproducibility of results, even with the same model and hardware. The particular software supply chain can be succinctly categorized into various runtimes, compilers, or implementations of high-level frameworks, such as PyTorch or Tensorflow. This highlights the need for careful version control and standardized compilation practices to mitigate potential discrepancies.
6.3 Hardware Supply Chains
A wide range of hardware is now available through various cloud platform providers, including GPUs from different vendors (NVIDIA, AMD, Intel) and specialized AI accelerators (Google TPUs, AWS Inferentia, Groq and Cerebras). It is also now popular for application developers to directly purchase model API access. For instance, emerging service providers like OpenRouter  [^2] https://openrouter.ai/ operate as marketplaces for LLM API. They offer a platform that aggregates LLM open model inference services from various service providers and even provide query blending, whereby queries are processed using a mix of cost-effective and expensive models. Typically, it is not fully disclosed to the users how these models are served. Limited information on the hardware used by such service providers, make it difficult for users to trust the results. Various hardware can introduce slight performance differences and without hardware platform accountability, API calls could be served from a provider implementing less secure or incompliant practices. The HSPI methods we have introduced a solution to check the potential hardware supply chain used by these service providers.
6.4 ML Governance
In the perfect HSPI scenario, our methods can help identify the exact components in a model’s software and hardware supply chains as mentioned in Section 6.2 and Section 6.3. This would allow users to produce and validate the software and hardware bill of materials, as suggested by recent research (Arora et al., 2022). By doing so, HSPI could become a key tool in enabling ML governance – a framework of standards and principles to ensure that an ML system operates responsibly upon deployment. The proposed HSPI problem formulation, including Schlögl et al.’s insights on EQCs, will help in shaping standards in different stages of an ML model’s life cycle from testing to deployment (Chandrasekaran et al., 2021).
Traceability and Accountability: In complex ML workflows, traceability of the exact software and hardware configurations used in model training and inference is essential for building trust and accountability. This would help in identifying the origin of specific model behavior differences, making reproduction of behavior easier. This traceability is not only crucial for debugging but also for audits, security, and mitigating risks from potential tampering or errors.
Establishing Industry Standards: HSPI can drive the adoption of industry-wide standards for documenting and verifying hardware and software configurations in ML development. Standardized reporting of hardware architectures, library versions, and compilers would enable researchers and practitioners to replicate results reliably. This would also help in establishing ML quality control norms.
Model Sharing and Deployment Reliability: Replicability through HSPI and newly established industry standards can further model sharing practices. Detailed information about the training and inference environment can reduce deployment issues related to hardware and software mismatches and can help users understand potential limitations or dependencies on specific hardware or software components. This will ultimately create an environment fostering confidence in using models developed by other people.
6.5 Robustness of HSPI
Several mitigation strategies can be employed to make the inference of hardware and software platforms significantly more difficult. These strategies focus on disrupting the patterns and subtle variations exploited by HSPI.

### v2

@@ 5.2 White-box Attacks
We find HSPI-BI always distinguishes between different number formats/quantization methods on the same GPU. Examples of border images can be found in Figure 4.  Figure 5 illustrates the kernel density estimation of various quantization methods where the shapes imply the obvious differences. When extending to real GPUs, we find the HSPI-BI generally works except for GPUs with very similar specs. For example, border images can be created between RTX8000 and H100 when running on identical CUDA version and PyTorch version, but not between RTX8000 and RTX2080Ti (see Table 8 in the appendix). This is because both RTX8000 and RTX2080Ti have the same Compute Capability of 7.5 and a very similar number of Tensor Cores around 550.
HSPI-LD can distinguish between all quantization levels for the models ResNet18, ResNet50, VGG16, EfficientNet, DenseNet, and MobileNet-V2 when using an SVM with a set of 10 images’ logits per sample. When comparing across GPUs, HSPI-LD are able to distinguish the A100 well from the RTX8000 and RTX2080Ti with an accuracy of $100\%$ (Figure 9 in the appendix shows the clear difference of logit bit distribution). Again, HSPI-LD fails to distinguish between the RTX8000 and RTX2080Ti, as calculations are in the same EQC. We discuss the limitations and robustness of HSPI in Appendix E and Appendix F.
@@
We find that HSPI-BI is not suitable for LLMs. With enough reruns and PGD iterations, HSPI-BI can generate border requests that lead to different responses across quantization methods. However, we find that the PGD process is unstable due to the discontinuity in its projection. Moreover, limited GPU memory constrains our ability to scale experiments to larger LLMs and bigger request batch sizes. This is because the forward pass of the target model is part of the computation graph for gradient descent, requiring caching the intermediate activations of LLMs in memory for backward propagation.
@@
Figure 6: The difference of bit distribution between RTXA6000 and A100 (white-box HSPI-LD). We send the same 256 queries to QWen-2.5-3B deployed on RTXA6000 and A100 respectively and compare the bit distribution of FP32 log probabilities. Tokens and logits are sampled in the plot but the difference is still obvious. $\mathrm{t_{i}l_{j}}$ denotes the log probability of $i$-th token’s $j$-th logit.
On the contrary, HSPI-LD achieves remarkable success in both quantization experiments and real GPU experiments. For quantized LLaMA-3.1-8B, Phi-3.5-mini, and QWen-2.5-3B deployed in FP16, BF16, INT8-FD, and HQQ-4bit, an accuracy of over 99.5% for each model is achieved (See Table 9 in the appendix). For real GPU experiments, we firstly perform three experiments in Table 1 to verify that HSPI-LD can distinguish between different GPUs, arithmetic modes, and kernel implementations. We then conduct a comprehensive experiment in Table 2, mixing all these factors. We treat this experiment as a classification task with 18 unique labels. Remarkably, HSPI-LD achieves an overall accuracy of 83.9% (random guess accuracy = 5.6%). We visualize the evident difference in logit bit distribution between RTXA6000 and A100 in Figure 6.
@@ 5.3 Black-box Attacks
(1) If border images are trained in batches, a larger batch size will improve the transferability across models. Figure 7(a) sweeps the batch size of border images trained on MobileNet-v3-Small and plots the resultant transferability. With the increasing batch sizes, the transferability is enhanced. However, we cannot further increase the batch size after 256 due to limited GPU memory.
(2) Overfitting happens if the border images are trained with too many iterations. Figure 7(b) presents the transferability of the same group of border images trained with various number of iterations. The decreasing transferability at 800 and 1200 iterations indicates that the border images overfit the training model.
(3) The border images trained on a more complex model tend to have better transferability. For example, in  Table 3, the border images trained on MobileNet-v2 has a higher F1-Score than those trained on ResNet18 and VGG16. However, again, more complex models require more GPU memory during training, which leads to limited scalability of HSPI-BI.
For HSPI-LD, we compare how well an SVM trained on one model performs on all other models in  Table 3. We observe that the transferability in detecting BF16 and FP16 is particularly high, while for other quantization levels such as FP32 and MXINT8 the transferability is very low. Furthermore, ResNet18, followed by DenseNet perform best as models with high transferability to other models.
@@
Here we summarize the promising approaches and challenges in Table 5 based on our experiments and observations. Though we achieve three to five times higher accuracy than random guess in the black-box setup, more effective solutions still need to be explored.
@@
Table 5: Promising approaches () and challenges (!) of HSPI.
Vision Models Language Models Setup White-box Black-box White-box Black-box HSPI-BI ! ! ! HSPI-LD !
@@ 6 Discussion
6.1 Tests on LLM Serving Frameworks
LLM serving frameworks like vLLM (Kwon et al., 2023) and SGLang (Zheng et al., 2024) facilitate the deployment of LLMs in production environments. In Section 5, we run the experiments with a naive setup. To verify HSPI-LD in production environments, we launch SGLang on cloud GPUs, send queries and collect responses through OpenAI API, and perform HSPI-LD. Though SGLang enables a series of optimizations, such as RadixAttention and cache-aware scheduling policy, HSPI-LD successfully recognizes different GPUs combined with various data types, partitioning strategies, and kernel implementations. Table 6 presents the results where the over 99% accuracy highlights the effectiveness in a production environment. In Appendix D, we include two more result tables on large models served on devices & software stack from various vendors like AMD, NVIDIA, and Amazon, highlighting that HSPI-LD is applicable to a more complex setup.
Table 6: White-box experiments of SGLang using cloud GPU service. We perform white-box HSPI-LD on a setup close to production environments where LLM serving framework, SGL, is deployed on a multi-GPU cloud server. We send queries and collect responses through OpenAI API.
GPUs Ariths. Kernels Sharding ClassIdx Acc. F1. L40 FP16 FlashInfer DP2, TP2 1 0.961 0.972 DP4, TP1 2 1 0.992 Triton DP2, TP2 3 0.992 0.9988 DP4, TP1 4 1 0.992 BF16 FlashInfer DP2, TP2 5 1 1 DP4, TP1 6 1 1 Triton DP2, TP2 7 1 1 DP4, TP1 8 1 1 A40 FP16 FlashInfer DP2, TP2 9 1 0.973 DP4, TP1 10 0.992 0.988 Triton DP2, TP2 11 0.961 0.953 DP4, TP1 12 0.953 0.976 BF16 FlashInfer DP2, TP2 13 1 1 DP4, TP1 14 1 1 Triton DP2, TP2 15 1 1 DP4, TP1 16 1 1 Average 0.991 0.991
6.2 Scalability and Cost Implications of HSPI
One may question the potentially high cost of HSPI due to the large number of possible combinations of hardware and software. As shown in previous experiments, the number of classes is the product of possible hardware platforms and software dependencies, e.g., ||||{L40, A40}x{FP16, BF16}x{FlashInfer, Triton}x{DP2TP2, DP4TP1}|=16|=16| = 16 classes in Table 6. If we decompose this classification task into binary classification problems, $\frac{N(N-1)}{2}=120$ classifiers need to be trained. However, we emphasize that one can reduce the number of classes by “merging labels”. By “merging labels” we mean merging some unimportant labels (don’t cares) into a new label. Take Table 6 as an example, if we label the eight classes {L40}x{FP16, BF16}x{FlashInfer, Triton}x{DP2TP2, DP4TP1} into a new label “L40” and the rest into “A40”, we can train a two-class classifier with a high F1 score of 98.5% (See Table 12 in the appendix). Another factor determining the cost of HSPI is the expense of LLM API access for collecting training samples. As detailed in Appendix B, we need around 60k tokens for each class, which is acceptable considering LLM API usually costs less than \$0.05 per 1k tokens (Cortenix & Bruner, 2024).
6.3 Limitations and Robustness
HSPI faces several practical limitations. Some configurations remain indistinguishable due to leaving computations in the same EQC - for instance, different CUDA versions on identical hardware, or similar GPUs (RTX8000 and RTX2080Ti) with matching compute capabilities. Moreover, batch size variations significantly affect logit outputs, requiring separate analysis for each common batch size. We discuss these limitations and possible strategies against the use of HSPI such as random bit flips in logits or noise injection in input data in Appendices E and F.
6.4 Software and Hardware Supply Chains
Software supply chain variations across runtimes, compilers, and high-level frameworks, such as PyTorch or Tensorflow, can affect result reproducibility even with identical models and hardware. This highlights the need for careful version control and standardized compilation practices to mitigate potential discrepancies on the software side. Meanwhile, hardware options have proliferated through cloud providers, ranging from various GPU vendors to specialized AI accelerators (Google TPUs, AWS Inferentia, Groq and Cerebras). The rise of model API marketplaces like OpenRouter [^2] https://openrouter.ai/ has introduced additional complexity, as these services often obscure their underlying hardware configurations. This lack of transparency in hardware platforms makes it difficult to verify security compliance and performance consistency. HSPI methods offer a potential solution for validating these hardware supply chains.
6.5 ML Governance
HSPI could significantly advance ML governance by enabling precise identification of hardware and software components in ML supply chains. This capability would strengthen three key areas: (1) traceability and accountability through detailed configuration tracking, (2) establishment of industry-wide standards for hardware and software documentation, and (3) improved reliability in model sharing and deployment. We will elaborate on these governance and broader impacts in our impact statement.
7 Conclusion
In this paper, we introduce Hardware and Software Platform Inference (HSPI) – a method for identifying the underlying hardware and software stack solely based on the input-output behavior of machine learning models. Our approach leverages inherent quantization limitations and the arithmetic operation orders of different GPU architectures to distinguish between various GPU types and software stacks. Our findings demonstrate the feasibility of inferring this information from black-box models, underscoring its implications for ensuring transparency and accountability in the LLM market. We will open-source our logit dataset and codes once the paper is accepted.
Impact Statement
HSPI’s potential applications extend beyond hardware identification to several aspects of ML system governance and deployment. Building on our earlier discussion of traceability, standardization, and reliability benefits, this section examines the broader technical and practical implications of HSPI. We consider its effects on security auditing practices, deployment verification protocols, and potential impacts on hardware provisioning strategies.
In the perfect HSPI scenario, our methods can help identify the exact components in a model’s software and hardware supply chains as mentioned in Section 6.4. This would allow users to produce and validate the software and hardware bill of materials, as suggested by recent research (Arora et al., 2022). By doing so, HSPI could become a key tool in enabling ML governance – a framework of standards and principles to ensure that an ML system operates responsibly upon deployment. The proposed HSPI problem formulation, including Schlögl et al.’s insights on EQCs, will help in shaping standards in different stages of an ML model’s life cycle from testing to deployment (Chandrasekaran et al., 2021).
Traceability and Accountability
In complex ML workflows, traceability of the exact software and hardware configurations used in model training and inference is essential for building trust and accountability. This would help in identifying the origin of specific model behavior differences, making reproduction of behavior easier. This traceability is not only crucial for debugging but also for audits, security, and mitigating risks from potential tampering or errors.
Establishing Industry Standards
HSPI can drive the adoption of industry-wide standards for documenting and verifying hardware and software configurations in ML development. Standardized reporting of hardware architectures, library versions, and compilers would enable researchers and practitioners to replicate results reliably. This would also help in establishing ML quality control norms.
Model Sharing and Deployment Reliability
Replicability through HSPI and newly established industry standards can further model sharing practices. Detailed information about the training and inference environment can reduce deployment issues related to hardware and software mismatches and can help users understand potential limitations or dependencies on specific hardware or software components. This will ultimately create an environment fostering confidence in using models developed by other people.
Appendix A Additional Related Work
Side-channel have been exploited to infer information about the model and hardware For example, Hua et al. (2018) extracts information about CNN models, e.g, whether the model is AlexNet or SqueezeNet, from an FPGA-based CNN accelerator, by analyzing memory access patterns. Recent work Gongye et al. (2024) predicts information about the encrypted IPs on an FPGA-based DNN accelerator. They also recover characteristics about model architectures and extract model params. Our work aims to predict information about the SW-HW stack used for serving large deep learning models, including compiler, data types, GPU arch, parallelism strategy, etc.
Our work is different from these two papers in terms of goals, threat models (methods), and generalizability.  Hua et al. (2018) and Gongye et al. (2024) assume side channel information about the HW like off-chip memory access patterns, electromagnetic traces, schematics is accessible, while we assume the SW-HW platform serves DNNs in the cloud and we only have access to the request and responses via the service provider’s API. These two works test their methods on a specific FPGA accelerator, while We tried our best to test the generalizability, across model families, data types, GPU archs, parallelism strategies, etc.
Appendix B Detailed Experiment Setup
We elaborate the detailed experiment setup in this sections. For the version information of software, please refer to our source codes.
B.1 Vision Model
Number formats and GPU Specs
We perform the experiments on standard floating-point formats (FP32, FP16, BF16) and low-precision formats first, then extend to actual GPUs. The low-precision formats include MXINT8 (Darvish Rouhani et al., 2020), FP8-E3M4 (noted as FP8-E3) (Shen et al., 2024), FP8-E4M3 (noted as FP8-E4), and dynamic INT8 (noted as INT8) (Kim et al., 2021). For actual GPUs, we include NVIDIA H100, A100, GeForce RTX2080 Ti, and Quadro RTX8000. Not all GPUs were used for all experiments due to different server locations and the difficulty of connecting all of them for the attack generation phase.
Border Images
Fine-tuning Hyper-parameters of Vision Models
The checkpoints of vision models in Section 5 are downloaded from torchvision 333 https://pytorch.org/vision/stable/models.html#classification. Before HSPI experiments, we fine-tune these models on CIFAR10 to ensure they achieve reasonable accuracy on CIFAR10 test set. Specifically, we use SGD optimizer and linear learning rate scheduler with initial learning rate = 1e-3. The fine-tuning batch size is 128 and we fine tune all the models for 3 epochs.
Accuracy of Quantized Models
We show in Table 7 that all low-precision formats have negligible accuracy loss compared to the fine-tuned model. This serves as a sanity check before the HSPI experiments in case the quantization breaks the models.
Table 7: Accuracy of quantized vision models on CIFAR10.
B.2 Language Models
Quantization methods and GPUs
Similar to vision models, we first run experiments of differentiating low-precision formats, then extend to actual GPUs. Since LLM quantization is more challenging than vision models, besides FP16 and BF16, we adopt two quantization methods that have been proven effective and integrated into HuggingFace: fine-grained dynamic INT8 quantization (Torchao, 2024), noted as INT8-FD, and half-quadratic quantization (Badri & Shaji, 2023), noted as HQQ-4bit. For actual GPUs, we consider NVIDIA A100, L40S, RTX A6000, and GeForce RTX3090.
Prompts and logit datasets
For HSPI-LD, we guide language models to generate random tokens with the adapted official chat template of each model.  Figure 8 shows an example of the template of QWen-2.5, where {num_random_words} and {random_words} are parameters for generating queries. {random_words} is $n$ leading random words sampled from the “popular” corpora of Natural Language Toolkit [^4] Natural Language Toolkit: https://www.nltk.org/index.html, thus leaving ({num_random_words} - $n$) tokens for the LLM to generate. Throughout our experiments, we find {num_random_words} = 16 and $n$ = 3 is enough for the white-box setup, while for the black-box setup, we adapt {num_random_words} = 16 and $n$ = 8.
Figure 8: The prompt template of QWen-2.5 used by HSPI-LD in Section 4.3.
We generate 256 queries for each class, and sample logits in Tables 9, 1 and 4 because of varying vocabulary sizes of LLMs. Specifically, we sample the top-256 logits of each generated token, and we flatten the logits of 4 responses for a training sample of SVM classifier.
Appendix C Additional Vision Model Results
Table 8 shows that HSPI-BI can differentiate RTX8000 and A100 but fails to differentiate RTX8000 and RTX2080Ti because they two GPUs have very similar specs. Figure 9 shows the clear difference in logit bit distribution between RTX8000 and NVIDIA A100, which explains the success of HSPI-LD.
Table 8: Table showing success in creating border images for different GPUs in white-box with an inference batch size of 1.
GPUs FP32 BF16 FP16 MXINT8 FP8-E3 FP8-E4 INT8 RTX8000 vs A100 RTX8000 vs 2080Ti
Appendix D Additional LLM Results
D.1 White-box HSPI-LD of quantized LLMs
Table 9 shows the results of HSPI-LD on quantization methods for LLMs with an average accuracy of 0.995.
D.2 HSPI-LD on Various GPU Systems
In Table 10, we distributed a large LLM, Llama-3.1-70B, across a larger multiple high-end GPU systems using SGLang. HSPI-LD still achieves high accuracy. We also test HSPI-LD on devices from different vendors, including NVIDIA, AMD, and Amazon in Table 11. Different devices are distinguished by their hardware architecture, software stack, and data types.
Vendor Hardware (arch) N devices DP,TP DType Acc F1. AMD MI300X (CNDA3) 2 DP1TP2 BF16 1.000 1.000 FP16 0.961 0.957 NVIDIA H100 (Hooper) 4 DP2TP2 BF16 1.000 1.000 FP16 0.930 0.941 L40 (Ada) 8 DP1TP8 BF16 1.000 1.000 FP16 0.969 0.958 A40 (Ampere) 8 DP1TP8 BF16 1.000 1.000 FP16 0.922 0.925 Avg 0.973 0.973
Table 10: Whitebox HSPI-LD on various real GPUs.
Vendor Hardware Software DType Accuracy F1. Amazon Inferentia NKI BF16 0.968 0.924 FP16 0.988 0.908 AMD MI300X ROCm BF16 1.000 1.000 FP16 0.906 0.913 NVIDIA H100 CUDA BF16 1.000 1.000 FP16 0.922 0.915 Avg 0.964 0.943
Table 11: White-box HSPI-LD on devices from various vendors.
D.3 Label merging
With the same samples in Table 6, we merge the samples of the same GPU but different software dependencies into a new label and train a new classifier. Surprisingly, the classifier still achieves a high accuracy, implying that label merging can effectively reduce the number of classes and improve scalability of HSPI-LD.
Table 12: White-box experiments of Table 6 with merged labels. We label the samples with the same GPU but various software as the same class and train a new classifer. The classifier still achieves high accuracy and F1 score, implying label merging is an effective way to reduce the number of labels and to improve the scalability of HSPI-LD.
GPUs Accuracy F1 Score A40x{FP16, BF16}x{FlashInfer, Triton}x{DP2TP2, DP4TP2} 0.990 0.985 L40x{FP16, BF16}x{FlashInfer, Triton}x{DP2TP2, DP4TP2} 0.979 0.985 Average 0.985 0.985
Appendix E Limitations and Black-box Failure Cases
@@
Appendix F Robustness of HSPI